OpenStudy (anonymous):

The linear approximation at x = 0 to f(x) = \sqrt { 8 + 2 x } is y =

OpenStudy (anonymous):

\[\sqrt{8+2x}\]

OpenStudy (coldplay):

x−−−−− Its supposed to be the square root sign \[\sqrt{8+2x}\]

OpenStudy (anonymous):

yes

OpenStudy (coldplay):

Got it.

OpenStudy (anonymous):

how

OpenStudy (anonymous):

Can someone help me with my question please http://openstudy.com/study#/updates/545aecf7e4b017f3f584ed4b

OpenStudy (freckles):

Use the same formula from last time: \[y=f(0)+f'(0)(x-0)\]

OpenStudy (anonymous):

so its 0?

OpenStudy (freckles):

what's 0?

OpenStudy (freckles):

\[f'(x)=?\]

OpenStudy (anonymous):

i mean sqrt of 8

OpenStudy (freckles):

well f(0) is sqrt(8)

OpenStudy (anonymous):

so the derivative of that is 2?

OpenStudy (coldplay):

Yes.

OpenStudy (freckles):

\[y=f(0)+f'(0)(x-0) \\ y=\sqrt{8}+f'(0)(x-0) \\ y=\sqrt{8}+f'(0)x\] The derivative of sqrt(8+2x) is not 2.

OpenStudy (coldplay):

Oh wait no its not 2

OpenStudy (camper4834):

\[(8+2x)^\frac{ 1 }{ 2 }\] do this :)

OpenStudy (anonymous):

its (2x+8)^-1/2

OpenStudy (camper4834):

no manhani that negative exponent would make it a recipricol

OpenStudy (camper4834):

wait what?

OpenStudy (anonymous):

thats the derivative of the equation tho

OpenStudy (camper4834):

XD yes yes sorry

OpenStudy (anonymous):

but when i plug in sqrt 8 + the derivative its saying its the wrong answer

OpenStudy (camper4834):

+ the derivative TIMES x

OpenStudy (camper4834):

did you forget the final x?

OpenStudy (anonymous):

2.82+x((2x+8)^(-1/2)) its not going through

OpenStudy (camper4834):

which site are you using, sometimes they want it simplified a certain way

OpenStudy (anonymous):

webwork its for my homework

OpenStudy (perl):

The equation of the tangent line to f(x) at x = a is L(x) = f(a) + f '(a) * ( x - a )

OpenStudy (camper4834):

\[\sqrt{8}+\frac{ x }{ \sqrt{8+2x} }\]

OpenStudy (camper4834):

y = what is above

OpenStudy (camper4834):

try that and tell me what happens

OpenStudy (anonymous):

incorrect

OpenStudy (camper4834):

double check the equation you gave us then

OpenStudy (anonymous):

The linear approximation at x = 0 to f(x) = \sqrt { 8 + 2 x } is y =

OpenStudy (anonymous):

did you multiply x?

OpenStudy (camper4834):

im stupid.... PLUG 0 INTO OUR DERIVATIVE!!

OpenStudy (camper4834):

(2*0+8)^-1/2

OpenStudy (anonymous):

i finally got it

OpenStudy (anonymous):

Thank you for your time!!

OpenStudy (camper4834):

:)